OpenStudy (anonymous):

Given that f(x)=x^3 find all values of x in the interval (-1, 1) that satisfy the mean value theorem.

OpenStudy (anonymous):

is this really how the question reads?

OpenStudy (anonymous):

points don't "satisfy the mean value theorem" a function does

OpenStudy (anonymous):

in any case we can compute \(\frac{f(1)-f(-1)}{1+1}=\frac{2}{2}=1\) and solve \[f'(x)=3x^2=1\] for \(x\)

OpenStudy (anonymous):

divide by 3, take the square root, don't forget the \(\pm\)

OpenStudy (anonymous):

Thank you.